[figure: 9.7 cm]

GIANT MAGNETORESISTANCE IN MELT-SPUN Co-Cu MODIFIED BY RARE EARTHS

B. Idzikowski, U.K. Rößler^{a}, A. Handstein^{a}, K. Nenkov^{a}^{∗}, K.-H. Müller^{a}

Institute of Molecular Physics, Polish Academy of Sciences, Poznań, Poland aInstitut für Festkörper- und Werkstofforschung Dresden, Germany

We present results on the modification of the giant magnetoresistance effect (GMR) in granular Co–Cu alloys. The samples of the composition RE_{x}Co_{20}Cu_{80−x}, RE = Sm, Gd or Dy (x=0, 1 and 3) were produced by melt–spinning and subsequent annealing. The magnetoresistance MR = [R(H) − R(0)]/R(0) is negative as in granular Co-Cu, but its values depend on the kind of rare earth element added and on changes in the microstructure induced by thermal treatment. The highest MR was achieved upon annealing at 450–480^{o}C for 1h. An addition of 1 at.% rare earth metal yields a considerable increase of MR compared to Co_{20}Cu_{80}, whereas an addition of 3 at.% causes a decrease of this value. E.g., by addition of 1 at.% RE=Dy to Co_{20}Cu_{80} and optimum thermal treatment of the alloys, the value of MR measured at 10 K increases from 15% to 28%. The MR ratio measured at 10 K for a ribbon with RE=Dy and x=1 does not reach saturation even in fields up to 16 T. The effects can be attributed to the presence of hexagonal Dy(Co,Cu)_{5} clusters, which was ascertained from X–ray diffraction data, and/or cubic DyCu_{5} besides fcc Co grains and diluted Co or RE atoms in the Cu matrix.

∗On leave from International Laboratory of High Magnetic Fields, Wrocław, and ISSP, BAS Sofia, Bulgaria.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : B. Idzikowski

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences M. Smoluchowskiego 17, 60-179 Poznań, Poland Fax: (48-61) 8684 -524

Email address : Bogdan.Idzikowski@ifmpan.poznan.pl